# DistillNeRF: Perceiving 3D Scenes from Single-Glance Images by Distilling Neural Fields and Foundation Model Features

### v1

@@ 3.2 Self-supervised Training with Distillation
Distillation from Offline NeRFs. While our model can be trained by simply reconstructing RGB images, it remains challenging to learn scene geometry from only single time-step camera image inputs. The challenge is especially pronounced with typical autonomous vehicle setups where mounted cameras are facing outwards and have limited view overlap, making multi-view reconstruction degrade to the monocular setting and aggravating depth ambiguity. A natural idea is to use images from multiple time steps to encourage consistent depth prediction. However, driving scenes typically contain many dynamic objects that move between time steps, introducing noise to the reconstruction objective. Instead, we propose to leverage the high-quality geometry of per-scene optimized NeRFs that aggregate information from a full sensor stream. Specifically, we use EmerNeRF [14], a recent NeRF approach that handles dynamic objects by decomposing the scene into static and dynamic fields in a self-supervised manner. We propose two different ways to distill knowledge from per-scene optimized NeRFs, which together construct $L_{NeRF}$, a distillation loss from offline NeRFs:
@@
Distillation from Foundation Models. In addition to RGB and depth prediction, can we learn 3D representations that enable a wider range of downstream applications? Witnessing the rise of vision foundation models with generalized capabilities across various vision tasks, we propose to distill 2D foundation model features, such as CLIP [15] and DINOv2 [16], into our 3D scene representation model. We achieve this by simply introducing an additional MLP to our rendered 2D feature images, and train the model to reconstruct the foundation model feature images with an L1 loss $L_{found}$. We demonstrate early attempts at utilizing these foundation models on open-vocabulary text query tasks shown in Fig 3, but we leave more comprehensive explorations to future work.
@@
where $L_{rgb}$ and $L_{depth}$ are rendering losses for RGB and depth; $L_{density}$ is a density entropy loss as in [14]; $L_{NeRF}$ and $L_{found}$ denote distillation losses from offline NeRFs and foundation models.
@@ 4 Experiments
We benchmark DistillNeRF against SOTA generalizable NeRFs, offline NeRFs, as well as comparable methods on the popular NuScenes dataset [51]. We evaluate rendering performance, i.e., scene reconstruction, novel view synthesis, and feature reconstruction (Table 1). We further evaluate the learned 3D geometry through depth estimation (Table 2), and 3D semantic occupancy prediction (Table 3). Ablations of DistillNeRF are also analyzed in each task. Qualitative results, including open vocabulary queries, are in Fig. 3, 4, and 5. Videos are online at https://distillnerf.github.io/. Implementation and training details are in the Appendix A.1.
@@
Dataset. The nuScenes dataset [51] contains 1000 driving scenes from different geographic areas, each scene capturing approx. 20 seconds of driving, resulting in approx 40000 frames in total. Scenes are captured via six cameras mounted on the vehicle heading in different directions along with point clouds from LiDAR. We use the default data split, 700 scenes for training, 150 scenes for validation. We adopt the resolution of input RGB images, rendered RGB, and rendered depth are 114×228, 114×228, and 64×114 respectively.
4.1 Rendering Images and Foundation Features
Setup. We evaluate models on previously unseen scenes from the validation set. For scene reconstruction, we compare the rendered images against GT images for the same time step. For novel-view synthesis, we render the novel-view image from the next timestep’s camera pose, and compare it against the next timestep’s GT image. We use standard metrics: peak signal-to-noise ratio (PSNR) and structural similarity index (SSIM). We compare with two SOTA generalizable NeRF methods, SelfOcc [22] and UniPAD [21], which have been successfully applied to outdoor driving scenes. We also compare with the SOTA per-scene optimized method, EmerNeRFs [14]. Since EmerNerfs are trained on all timesteps in the scene, we cannot evaluate them for novel views. Instead, for novel-view evaluation we adapt Single-Frame EmerNeRFs, each of which is trained only on a single timestep and then evaluated for the next timestep. Due to the prohibitive training cost of Single-Frame EmerNerf on all timesteps, for all methods we report mean metrics over only the second frame of each scene.
RGB Reconstruction and Novel-View Synthesis. In Table 1, we compare DistillNeRF with existing approaches on the image reconstruction and novel-view synthesis on the nuScenes validation set. The results show that our generalizable model is on par with the per-scene optimized NeRFs, and significantly outperforms SOTA generalizable methods, both for RGB reconstruction and novel-view synthesis. Reconstruction PSNR for our best model variant is close to per-scene optimized EmerNerfs (30.11 vs 30.88), and achieves even higher SSIM (0.917 vs 0.879). Similarly, our novel-view PSNR is close to Single-Frame EmerNerf (20.78 vs 20.95), while SSIM is slightly higher (0.590 vs 0.585). Compared prior SOTA generalizable methods, our model outperforms the best-performing method (SelfOcc) in PSNR by 45.6% and 14.0%, and in SSIM by 64.9% and 27.1%, for reconstruction and novel-view synthesis, respectively. Novel-view metrics are generally lower than reconstruction metrics. Note that our novel-view setting is challenging, as the vehicle can travel large distances (in 0.5s) between the input-view and novel-view camera poses, and capture elements that are invisible in the original camera pose. Further, dynamic objects that move between frames act as noise in our novel-view targets.
Qualitative results for reconstructions are in Fig. 4. EmerNerf and our approach are close to the ground truth, while UniPAD generates blurry reconstructions with scan patterns, and SelfOcc generates grayish images and struggles to reconstruct the scene precisely. Additional novel-view renderings are at https://distillnerf.github.io/.
Ablations. Ablation results in Table 1 and Fig. 5 indicate that depth distillation from offline NeRFs increases reconstruction and novel-view synthesis performance, while virtual camera distillation helps novel-view synthesis. The parameterized space slightly reduces the rendering metrics, but as shown in Fig 5, it is capable of generating unbounded depth.
@@ 4.1 Rendering Images and Foundation Features
Table 1: Reconstruction and novel-view synthesis on nuScenes validation set. DistillNeRF is on par with the per-scene optimized NeRFs, both in RGB and foundation feature rendering, and significantly outperforms SOTA generalizable NeRF methods. See Fig. 4 and Fig. 5 for qualitative results.
@@
Ours 28.01 0.872 19.12 0.501 - -
Ours (+ Depth Distillation) 30.11 0.917 20.27 0.567 18.69 18.48
Ours (+ Param Space) 28.42 0.879 20.06 0.565 - -
Ours (+ Virtual Cam Distillation) 28.72 0.880 20.78 0.590 - -
@@
Ours 0.248 3.090 6.096 0.312 0.704 0.885 0.947
Ours (+ Depth Distillation) 0.233 2.890 5.890 0.296 0.703 0.881 0.945
Ours (+ Virtual Cam Distillation) 0.223 1.776 5.461 0.293 0.763 0.903 0.961
@@
Ours 0.270 3.670 6.301 0.389 0.653 0.826 0.886
Ours (+ Depth Distillation) 0.235 3.008 5.859 0.311 0.726 0.890 0.942
Ours (+ Virtual Cam Distillation) 0.228 1.898 5.654 0.302 0.689 0.879 0.943
@@ 4.2 Depth Estimation
Setup. We evaluate depth up to 80m using common metrics (Abs Rel, Sq Rel, RMSE, RMSE log, and $\delta<t$) [52, 53, 54]. We use two different depth targets: Sparse LiDAR GT, the common evaluation setting using LiDAR point cloud to define depth, which is accurate but spare and has limited range (e.g. only  3m height); and Dense Depth GT, that uses EmerNerf to define dense depth targets with large range. We compare against the same baselines as for rendering. For UniPAD we increase the maximum range to 80m and retrain the model. For SelfOcc, we evaluate two model variants, SelfOcc* that supports depth prediction only (used in [22]), and SelfOcc that also supports rendering (thus more similar to our method). Same as before, we evaluate over the second frame of each scene.
@@
Ablations. Consistently with previous results, distillation from offline NeRFs also improve depth estimation (Table 2). Quantitatively (Fig. 5), without depth distillation, we see inconsistent depth predictions between low and high regions of the image. Without parameterized space, the model can only predict depth in a limited depth range, while with parameterized space we can generate reasonable unbounded depth.
@@
-0cm-0cm Method F-mIou mIoU G-IoU $\blacksquare$ others $\blacksquare$ barrier $\blacksquare$ bicycle $\blacksquare$ bus $\blacksquare$ car $\blacksquare$ cons. veh. $\blacksquare$ motorcycle $\blacksquare$ pedestrian $\blacksquare$ traffic cone $\blacksquare$ trailer $\blacksquare$ truck $\blacksquare$ drive. surf. $\blacksquare$ other flat $\blacksquare$ sidewalk $\blacksquare$ terrain $\blacksquare$ manmade $\blacksquare$ vegetation SimpleOcc [58] 3.68 - 7.99 - 0.67 1.18 3.21 7.63 1.02 0.26 1.80 0.26 1.07 2.81 40.44 - 18.30 17.01 13.42 10.84 OccNeRF [23] 5.33 - 10.81 - 0.83 0.82 5.13 12.49 3.50 0.23 3.10 1.84 0.52 3.90 52.62 - 20.81 24.75 18.45 13.19 SelfOcc (BEV) [22] 2.71 6.76 44.33 0.00 0.00 0.00 0.00 9.82 0.00 0.00 0.00 0.00 0.00 6.97 47.03 0.00 18.75 16.58 11.93 3.81 SelfOcc (TPV) [22] 4.14 9.30 45.01 0.00 0.15 0.66 5.46 12.54 0.00 0.80 2.10 0.00 0.00 8.25 55.49 0.00 26.30 26.54 14.22 5.60 Ours 3.48 4.63 13.41 0.02 0.77 1.41 5.77 6.33 1.56 1.32 4.38 3.3 0.47 4.34 20.14 0.00 8.36 8.44 4.76 7.37 Ours (+ Depth Distillation) 6.40 8.93 29.11 0.03 1.35 2.08 10.21 10.09 2.56 1.98 5.54 4.62 1.43 7.90 43.02 0.00 16.86 15.02 14.06 15.06

### v2

@@ 3.2 Self-supervised Training with Distillation
Distillation from Offline NeRFs. While our model can be trained by simply reconstructing RGB images, it remains challenging to learn scene geometry from only single time-step camera image inputs. The challenge is especially pronounced with typical autonomous vehicle setups where mounted cameras are facing outwards and have limited view overlap, making multi-view reconstruction degrade to the monocular setting and aggravating depth ambiguity. A natural idea is to use images from multiple time steps to encourage view overlaps. However, driving scenes typically contain many dynamic objects that move between time steps, introducing noise to the reconstruction objective. Instead, we propose to leverage the high-quality geometry of per-scene optimized NeRFs that aggregate information from a full sensor stream. Specifically, we use EmerNeRF [14], a recent NeRF approach that handles dynamic objects by decomposing the scene into static and dynamic fields in a self-supervised manner. We propose two different ways to distill knowledge from per-scene optimized NeRFs, which together construct $L_{NeRF}$, a distillation loss from offline NeRFs:
@@
Distillation from Foundation Models. In addition to RGB and depth prediction, can we learn 3D representations that contain rich semantics and enable a wider range of downstream tasks? Witnessing the rise of vision foundation models with generalized capabilities across various vision tasks, we propose to distill 2D foundation model features, such as CLIP [15] and DINOv2 [16], into our 3D scene representation model. We achieve this by simply introducing an additional MLP to our rendered 2D feature images, and train the model to reconstruct the foundation model feature images with an L1 loss $L_{found}$. We demonstrate early attempts at utilizing these foundation models on open-vocabulary text query tasks shown in Fig 3, but we leave more comprehensive explorations to future work.
@@
where $L_{rgb}$ and $L_{depth}$ are rendering losses for RGB and depth; $L_{density}$ is a density entropy loss encouraging clearer surfaces and structured density values as in [14]; $L_{NeRF}$ and $L_{found}$ denote distillation losses from offline NeRFs and foundation models. Please refer to Appendix A.5 for more details on the losses.
@@ 4 Experiments
We benchmark DistillNeRF against SOTA generalizable NeRFs, offline NeRFs, as well as comparable methods on the popular NuScenes dataset [52]. We first evaluate the rendering performance, i.e., scene reconstruction, novel view synthesis, and feature reconstruction (Table 1). We further evaluate the learned 3D geometry through depth estimation (Table 2), and 3D semantic occupancy prediction (Table 4). Ablations of DistillNeRF are also analyzed in each task and displayed in each table, additional ablations are available in Tab A.2, Tab 6, and Tab 7 of the Appendix. Qualitative results, including open vocabulary queries, are in Fig. 3, 6, and 7. Videos are online at https://distillnerf.github.io/. Implementation and training details are in the Appendix A.7.
@@
Dataset. The nuScenes dataset [52] contains 1000 driving scenes from different geographic areas, each scene capturing approx. 20 seconds of driving, resulting in approx 40000 frames in total. Scenes are captured via six cameras mounted on the vehicle heading in different directions along with point clouds from LiDAR. We use the default data split, 700 scenes for training, 150 scenes for validation. We adopt the resolution of input RGB images, rendered RGB, and rendered depth are 114×228, 114×228, and 64×114 respectively. We also evaluate the generalizability of our method on the Waymo NOTR dataset [14], a balanced and diverse benchmark derived from the Waymo Open Dataset [26]. NOTR features 120 unique, hand-picked driving sequences, split into 32 static, 32 dynamic, and 56 diverse scenes across seven challenging conditions. We adopt the resolution of input RGB images, rendered RGB, and rendered depth are 144×216, 144×216, and 72×108 respectively.
@@
Table 1: Reconstruction and novel-view synthesis on nuScenes validation set. DistillNeRF is on par with the per-scene optimized NeRFs, both in RGB and foundation feature rendering, and significantly outperforms SOTA generalizable NeRF methods. In the DistillNeRF variants, we denote ’Depth’ as the depth distillation from offline NeRFs, ’Param.’ as the parameterized space, and ’Virt.’ as the distillation from virtual cameras in offline NeRFs. See Fig. 6 and Fig. 7 for qualitative results.
@@
Depth    |    Param.    |     Virt. DistillNeRF Variants
                               28.01 0.872 19.12 0.501 - -
                               30.11 0.917 20.27 0.567 18.69 18.48
                               28.42 0.879 20.06 0.565 - -
                               28.72 0.880 20.78 0.590 - -
@@
Depth Distill  |  Virtual Distill DistillNeRF Variants
                    |           0.248 3.090 6.096 0.312 0.704 0.885 0.947
                    |           0.233 2.890 5.890 0.296 0.703 0.881 0.945
                    |           0.223 1.776 5.461 0.293 0.763 0.903 0.961
@@
Depth Distill  |  Virtual Distill DistillNeRF Variants
                    |           0.270 3.670 6.301 0.389 0.653 0.826 0.886
                    |           0.235 3.008 5.859 0.311 0.726 0.890 0.942
                    |           0.228 1.898 5.654 0.302 0.689 0.879 0.943
4.1 Rendering Images and Foundation Features
Setup. We evaluate our model on previously unseen scenes from the validation set. For scene reconstruction, we compare the rendered images against GT images for the same time step. For novel-view synthesis, we render the novel-view image from the next timestep’s camera pose, and compare it against the next timestep’s GT image. We use standard metrics: peak signal-to-noise ratio (PSNR) and structural similarity index (SSIM). We compare with two SOTA generalizable NeRF methods in driving scenes, SelfOcc [22] and UniPAD [21], which do not see the validation set during training, similar to ours. We also compare with the SOTA per-scene optimized method, EmerNeRFs [14], which is trained on the validation set. Since EmerNerfs are trained on all timesteps in the scene, we cannot evaluate them for novel views. Instead, for novel-view evaluation we adapt Single-Frame EmerNeRFs, each of which is trained only on a single timestep and then evaluated for the next timestep. Due to the prohibitive training cost of Single-Frame EmerNerf on all timesteps, for all methods we report mean metrics over only the second frame of each scene.
RGB Reconstruction and Novel-View Synthesis. In Table 1, we show the results the image reconstruction and novel-view synthesis on the nuScenes validation set. The results show that our generalizable model is on par with the per-scene optimized NeRFs, and significantly outperforms SOTA generalizable methods, both for RGB reconstruction and novel-view synthesis. Without per-scene optimization, reconstruction PSNR for our best model variant is close to per-scene optimized EmerNerfs (30.11 vs 30.88), and achieves even higher SSIM (0.917 vs 0.879). Similarly, our novel-view PSNR is close to Single-Frame EmerNerf (20.78 vs 20.95), while SSIM is slightly higher (0.590 vs 0.585). Compared prior SOTA generalizable methods, our model outperforms the best-performing method (SelfOcc) in PSNR by 45.6% and 14.0%, and in SSIM by 64.9% and 27.1%, for reconstruction and novel-view synthesis, respectively. Novel-view metrics are generally lower than reconstruction metrics. Note that our novel-view setting is challenging, as the vehicle can travel large distances (in 0.5s) between the input-view and novel-view camera poses, and capture elements that are invisible in the original camera pose. Further, dynamic objects that move between frames act as noise in our novel-view targets. Qualitative results are in Fig. 6. EmerNerf and our approach are close to the ground truth, while UniPAD generates blurry reconstructions with scan patterns, and SelfOcc generates grayish images and struggles to reconstruct the scene precisely.
Table 3: Trained on the nuScenes dataset, DistillNeRF shows strong generalizability to the unseen Waymo NOTR dataset.
Method PSNR SSIM
EmerNeRF 28.87 0.814
Zero-Shot Transfer 21.03 0.841
Zero-Shot Transfer + Recolor 24.85 0.867
Finetune 29.84 0.911
Generalization to Unseen Waymo NOTR Dataset As in Fig 4 and Tab 3, we evaluate the generalizability of DistillNeRF to unseen domains. Trained on the nuScenes dataset, our model demonstrates strong zero-shot transfer performance on the unseen Waymo NOTR dataset. This quality can be further enhanced by applying simple color alterations to account for camera-specific coloring discrepancies. After fine-tuning, our model surpasses the offline per-scene optimized EmerNeRF on the data, achieving higher PSNR (29.84 vs. 28.87) and SSIM (0.911 vs. 0.814). We use the full Waymo NOTR data for evaluation, and quote original EmerNeRF metrics.
Ablations. Ablation results in Table 1 and Fig. 7 indicate that depth distillation from offline NeRFs increases reconstruction and novel-view synthesis performance, while virtual camera distillation benefits novel-view synthesis. The parameterized space slightly reduces the rendering metrics, but as shown in Fig 7, it is capable of generating unbounded depth.
@@ 4.2 Depth Estimation
Setup. We evaluate depth up to 80m using common metrics (Abs Rel, Sq Rel, RMSE, RMSE log, and $\delta<t$) [53, 54, 55]. We use two different depth targets: Sparse LiDAR GT, the common evaluation setting using LiDAR point cloud as ground truth, which is accurate but spare and has limited range (e.g. only  3m height); and Dense Depth GT, that uses EmerNerf to define dense depth targets with large range. We compare against the same baselines as for rendering. For UniPAD we increase the maximum range to 80m and retrain the model. For SelfOcc, we evaluate two model variants, SelfOcc* that supports depth prediction only (used in [22]), and SelfOcc that also supports rendering (thus more similar to our method). Same as before, we evaluate over the second frame of each scene.
@@
Ablations. Consistently with previous results, distillation from offline NeRFs also improve depth estimation (Table 2). Quantitatively (Fig. 7), without depth distillation, we see inconsistent depth predictions between low and high regions of the image; without parameterized space, the model can only predict depth in a limited depth range, while with parameterized space we can generate reasonable unbounded depth.
@@
-0cm-0cm Method F-mIou mIoU G-IoU $\blacksquare$ others $\blacksquare$ barrier $\blacksquare$ bicycle $\blacksquare$ bus $\blacksquare$ car $\blacksquare$ cons. veh. $\blacksquare$ motorcycle $\blacksquare$ pedestrian $\blacksquare$ traffic cone $\blacksquare$ trailer $\blacksquare$ truck $\blacksquare$ drive. surf. $\blacksquare$ other flat $\blacksquare$ sidewalk $\blacksquare$ terrain $\blacksquare$ manmade $\blacksquare$ vegetation SimpleOcc [59] 3.68 - 7.99 - 0.67 1.18 3.21 7.63 1.02 0.26 1.80 0.26 1.07 2.81 40.44 - 18.30 17.01 13.42 10.84 OccNeRF [23] 5.33 - 10.81 - 0.83 0.82 5.13 12.49 3.50 0.23 3.10 1.84 0.52 3.90 52.62 - 20.81 24.75 18.45 13.19 SelfOcc (BEV) [22] 2.71 6.76 44.33 0.00 0.00 0.00 0.00 9.82 0.00 0.00 0.00 0.00 0.00 6.97 47.03 0.00 18.75 16.58 11.93 3.81 SelfOcc (TPV) [22] 4.14 9.30 45.01 0.00 0.15 0.66 5.46 12.54 0.00 0.80 2.10 0.00 0.00 8.25 55.49 0.00 26.30 26.54 14.22 5.60        Depth Distill DistillNeRF Variants                3.48 4.63 13.41 0.02 0.77 1.41 5.77 6.33 1.56 1.32 4.38 3.3 0.47 4.34 20.14 0.00 8.36 8.44 4.76 7.37                6.40 8.93 29.11 0.03 1.35 2.08 10.21 10.09 2.56 1.98 5.54 4.62 1.43 7.90 43.02 0.00 16.86 15.02 14.06 15.06